OpenStudy (anonymous):

Which of the following quadratic expressions factors more than one time? Hint: Factor each one. Then go back and try to factor them again Answer a. 5x²+4x-12 b. 25m²-10m+1 c. 36n-9n² d. 5n²+50n+80

OpenStudy (anonymous):

- not sure exactly what you mean but for d : 5n²+50n+80 = 5(n^2 + 10n + 16) = 5 (n+8)(n+2)